OpenStudy (anonymous):

Calculate the molarity of 1.16 L of a solution containing 2.06 g of dissolved KBr.

OpenStudy (anonymous):

Molecular Mass of KBr= 39+80=119g. Moles Of KBr=weight/Molecular weight=2.06/119=0.017. Molarity= Moles/Volume= 0.017/1.16=0.014